OpenStudy (anonymous):

Use a graph of f or some other method to determine what, if any, value to assign to f(a) to make f continuous at x = a. f(x)= (x) / (7x^2-x); a= 0

OpenStudy (anonymous):

Just find the limit as x approaches zero \[\lim_{x \rightarrow a} f(x) = x/(7x^2 - x)\] first lets factor out an x from the bottom and rewrite it as: f(x) = x/x(7x - 1) now since you have an x on top and an x on bottom, you can just cancel them out: = 1/(7x - 1) now we plug in (a). Since (a) = 0, we can just plug in 0 = 1/(7(0) - 1) = 1/(0-1) = 1/-1 = -1 so \[\lim_{x \rightarrow a} f(x) = -1\] Since f(x) is undefined at 0, but the limit exists, we just need to define f(x) at zero to make it continuous. f(x) = -1 when x=-1